# Prompt Tuning Strikes Back: Customizing Foundation Models with Low-Rank Prompt Adaptation

### v1

@@ 5 Conclusion
Limitations and Future Work. The main limitation of LOPA is that its effectiveness as an alternative to LoRA was demonstrated on a set of benchmark tasks, but this may not hold for obscure real-life user tasks where LoRA or even full fine-tuning might be necessary. Further investigation into its performance on real-world tasks is part of our future work. In this work, we assumed the learned soft prompt to be prepended as a prefix to the input. Future research could explore the effects of positioning it as a suffix or randomly within the input. Finally, the foundation model combined with LOPA can be viewed as a Conditional Auto-Encoder, where soft-prompt vectors exist in a latent subspace rather than an offset subspace. This viewpoint raises intriguing questions, such as whether the observed performance improvements result from compressing knowledge from different instances and providing it as additional information. Investigating this alternate perspective could lead to further performance enhancements by developing more sophisticated auto-encoding systems.

### v2

@@ 5 Conclusion
Limitations and Future Work. The main limitation of LoPA is that its effectiveness as an alternative to LoRA was demonstrated on a set of benchmark tasks, but this may not hold for obscure real-life user tasks where LoRA or even full fine-tuning might be necessary. Further investigation into its performance on real-world tasks is part of our future work. In this work, we assumed the learned soft prompt to be prepended as a prefix to the input. Future research could explore the effects of positioning it as a suffix or randomly within the input. Finally, the foundation model combined with LoPA can be viewed as a Conditional Auto-Encoder, where soft prompt vectors exist in a latent subspace rather than an offset subspace. This viewpoint raises intriguing questions, such as whether the observed performance improvements result from inferring and compressing task-specific knowledge and providing it as additional information. Investigating this alternate perspective could lead to further performance enhancements by developing more sophisticated auto-encoding systems.
@@
6 Acknowledgments
This research was supported by the NSF under grant numbers CCF1918651, 2008240, 2131294, and 2212557, NIH CTSA award No. UL1TR003167, and US DOT Tier-1 UTC CYBER-CARE grant #69A3552348332.